OpenStudy (jewlzme17):

What is a double root and a triple root? I am graphing ana dI am suppose to have at least 5 touches to the x-axis with multiplicity of at least one double root and one triple root

OpenStudy (mathstudent55):

Multiplicity is the number of times the same root appears. Here is an example. Let's say you have equation \(x^2 + 3x + 2 = 0\) You factor it into \((x + 2)(x + 1) = 0\) and you get roots \(y = -2\) or \(x = -1\) Ok so far?

OpenStudy (mathstudent55):

Each of the above roots only appears once, so the multiplicity of each root is 1. Now let's say you have a 3rd-degree polynomial that factors into \((x + 2)(x + 2)(x + 1) = 0\) By applying the zero-product rule, you continue to solve it like this: x + 2 = 0 or x + 2 = 0 or x + 1 = 0 x = -2 or x = -2 or x = -1 Notice that the root x = -2 appears twice, so you can say the solutions are x = -2 with multiplicity 2 or x = -1

OpenStudy (mathstudent55):

A polynomial equation has as many solutions as the degree of the polynomial. For example, a 5th degree polynomial has 5 solutions. If the polynomial equation is \((x - 10)^5 = 0\) Then the solution is x = 10 with multiplicity 5, meaning there are 5 solutions, but they're all the same, x = 10.

OpenStudy (mathstudent55):

A double root is a root with multiplicity 2, and a triple root is a root with multiplicity 3.

OpenStudy (jewlzme17):

okay thank you! When graphing how does that apply? because it can only hit, let's say -2, once.

OpenStudy (amistre64):

what does a 5th degree poly look like when graphed?

OpenStudy (amistre64):

at least 5 touches .... so it can be greater than degree 5

OpenStudy (amistre64):

pick 2 values to be the triple and double roots, then pick 3 more values to hit the x axis whereever you please

OpenStudy (amistre64):

y=(x-a)^3 (x-b)^2 (x-c) (x-d) (x-e) it has 5 places it touches the x axis, and has multiplicity of roots required